Burke Tires just paid a dividend of D 0 = ₱ 1.32. Analysts expect the company's dividend to grow by 30%
this year, by 10% in Year 2, and at a constant rate of 5% in Year 3 and thereafter. The required return on this
low-risk stock is 9.00%. What is the best estimate of the stock’s current market value?
1
Expert's answer
2020-11-20T07:20:25-0500

"D_{0} = 1.32"

"D_{1}= 1.32 \times1.30 = 1.76, D_{2} = 1.76 \times1.10=1.936"

now "D_{3} =1.936\times1.05= 2.0328"

constant growth rate "=5\%"

required return "=9\%"

"P_{2} = \frac{D_{3}}{r-g} = \frac{2.0328}{0.09-0.05}=\$50.82"

"P_{0}=\frac{1.76}{1.09}+\frac{50.82}{1.09^{2}}= 1.6147 +42.7742 = \$44.39"